# Medical Spoken Named Entity Recognition

### v2

@@ 5 Conclusion
In this work, we present VietMed-NER - the first spoken NER dataset in the medical domain. Our dataset contains 18 entity types, including both conventional and newly defined entity types for real-world medical conversations. Furthermore, we present an incremental annotation technique based on gazeeteer list, called Recursive Greedy Mapping, to accelerate the annotation speed and remove inconsistency during annotating. Finally, we found that pre-trained multilingual models XLM-R_base and XLM-R_large outperformed all monolingual models on both reference text and ASR output, and in general, encoders perform better than seq2seqs for the NER task.
@@ 6 Limitations
Our annotation approach: Our annotation approach using the Recursive Greedy Mapping technique has some advantages over the fully manual approach. First, it allows annotators to not spend extra time tagging the entities that have been tagged in previous sentences. Second, it prevents that annotators miss entities that have been tagged in previous sentences, improving the consistency of the entire dataset. During our work, we experienced a faster annotation by using our approach compared to fully manual annotation. However, in the scope of this paper, we have not done extensive experiments to give a quantitative number of how much time has been saved and the method’s impact on annotation quality.
@@ Appendix A Annotation Guidelines
  • Include all jobs that might be both in medical fields and non-medical fields. For example: "khán thính giả" (audience), "kẻ bắt cóc" (kidnapper), "bệnh nhân" (patient), "người dân" (citizen), "chuyên gia" (expert).
@@
Appendix B Discussion about NEER
@@ B.5 Open questions on NEER
Appendix C Modified evaluation of SLUE toolkit
@@
We introduce a "dummy" token strategy to replace the actual named entities. This approach upholds the focus on the classification of entities rather than the extraction of verbatim phrases, which is suitable for cases where ASR errors might skew the recognition of entities in spoken transcripts.
@@
We compute two types of overall metrics: micro and macro averages. The micro average metrics aggregate the contributions of all classes to compute the average metric, while the macro average computes per-entity type metrics and averages them, without considering the frequency of each entity type. The micro average is therefore influenced by the class distribution and will be dominated by the performance on more frequent entity types. In contrast, macro averages treat all entity types equally, providing a measure of the system’s performance across different types of named entities, regardless of their frequency in the dataset.
Appendix D Sequence-to-sequence Training for NER Task
Following the approach proposed by Phan et al. (2021) and later adopted by ViT5 Phan et al. (2022), we formulated the token classification as a sequence-to-sequence task by training the models to generate tags of labels before and after an entity token. In cases where the models fail to follow the mentioned format for an entity token, we use an “exception” tag, which will be later ignored during metric calculation, as the label.
Appendix E Details of Training Hyperparameters
We used HuggingFace Wolf et al. (2019) for fine-tuning pre-trained models for the NER task. We only employed word-level settings for training NER models. All our NER experiments were done using the default hyperparameters by HuggingFace.
Appendix F NER Results by Entity Types
Table LABEL:appx_NER_results_byLabel shows the results of NER on reference text by entity types using various pre-trained language models. Table LABEL:appx_spk_NER_results_byLabel_ shows the results of NER on ASR output by entity types using various pre-trained language models and ASR models.
Figure 1 shows the scatter plot of NER results on reference text by entity types using various pre-trained language models, created by Table LABEL:appx_NER_results_byLabel. Figure 2 shows the scatter plot of NER results on ASR output by entity types using various pre-trained language models, created by Table LABEL:appx_spk_NER_results_byLabel_.
Table 5: NER results by entity types (in percent) on reference text of test set using various pre-trained language models. Metrics shown are Precision, Recall, and overall micro/macro F1 score. Evaluation toolkits used are: seqeval, SLUE and our modified SLUE.

### v3

@@ 7 Conclusion
In this work, we present VietMed-NER - the first spoken NER dataset in the medical domain. Our dataset contains 18 entity types, including both conventional and newly defined entity types for real-world medical conversations. Our results show that pre-trained multilingual models typically outperform monolingual models on both reference text and ASR output if the multilingual models are sufficiently large to learn multilingual representations. Additionally, encoders generally demonstrate better performance than seq2seq models in the NER task. Finally, while pre-trained audio data impacts ASR output, it does not significantly impact NER performance in the cascaded setting.
@@ 8 Limitations
Our annotation approach: Our annotation approach has some advantages over the fully manual approach. First, it allows annotators to not spend extra time tagging the entities that have been tagged in previous sentences. Second, it prevents that annotators miss entities that have been tagged in previous sentences, improving the consistency of the entire dataset. During our work, we experienced a faster annotation by using our approach compared to fully manual annotation. However, in the scope of this paper, we have not done extensive experiments to give a quantitative number of how much time has been saved and the method’s impact on annotation quality.
Evaluation metrics for medical terms: ASR system performance is commonly evaluated using WER, which quantifies the ratio of word insertion, substitution, and deletion errors in a transcript relative to the total number of spoken words. However, various spoken language understanding tasks, such as spoken NER, rely on accurately identifying key terms within transcripts. In medical ASR, it is critical to account for the disproportionate importance of medical terms in doctor-patient interactions, as they hold significantly more weight than general vocabulary, as discussed in Section B in the Appendix. We believe that other domain-specific spoken NER tasks follow a similar pattern. Consequently, future comprehensive investigations into evaluation metrics are needed to determine the most appropriate metric for spoken NER in the medical and other domains.
@@ Appendix A Annotation Guidelines
  • We adhere to the conventional approach of annotating overlapping NE components as a whole, utilizing the BIO encoding scheme. In recent years, research on overlapping and discontinuous NER has introduced alternative annotation frameworks to solve NE overlapping, such as BILOU encoding, which represents "Beginning, Inside, and Last tokens of multi-token chunks, Unit-length chunks, and Outside" Straková et al. (2019); Duan et al. (2017). Another approach by Li et al. (2021b) introduces a novel span-based model capable of jointly recognizing both overlapping and discontinuous entities. The model operates in two primary stages. First, entity fragments are identified by systematically traversing all possible text spans, enabling the detection of overlapping entities. Second, a relation classification step determines whether a given pair of entity fragments exhibits an overlapping or successive relationship. This approach facilitates the recognition of discontinuous entities while simultaneously verifying overlapping entities.
@@
  • Include all jobs that might be both in medical fields and non-medical fields. For example: "khán thính giả" (audience), "bệnh nhân" (patient), "người dân" (citizen), "chuyên gia" (expert).
@@
Appendix B Discussion about Named-Entity-Error-Rate (NEER)
@@ B.5 Open questions on NEER
Appendix C Possible Applications
In the context of advanced in-context learning capabilities of LLMs and human-level text-to-speech technologies, the dataset, which preserves entity positional labels throughout the translation process, is applicable not only to Vietnamese but also to other languages.
To facilitate dataset translation using LLMs, we employed a concise, brief contextual instruction via GPT-3.5 Turbo, as illustrated below.
In this example, we utilized a straightforward one-shot in-context learning approach, which successfully produced a high-quality translation while preserving all entity positional labels. We hypothesize that employing many-shot learning, extending the context length of instructions, and more advanced LLMs (e.g. GPT-4, GPT-4o) could yield even more enhanced results.
Recent research in text-to-speech (TTS) technology has achieved human-level performance in aspects such as naturalness and emotional expressiveness Tan et al. (2024); Li et al. (2023); Chen et al. (2023). Consequently, the application of TTS to translated NER transcripts could generate realistic audio data, facilitating further research in additional languages.
gpt-3.5-turbo Instruction: Given the text: "thì cũng giống như ba má mình đã từng bị béo phì rồi bị bệnh này bệnh kia những người thân quen của mình mình biết bị" Its coressponding named entity position is: [ "0", "0", "0", "0", "B-GENDER", "B-GENDER", "0", "0", "0", "0", "B-DISEASESYMTOM", "I-DISEASESYMTOM", "0", "B-DISEASESYMTOM", "I-DISEASESYMTOM", "0", "0", "0", "0", "0", "0", "0", "0", "0", "0", "0", "0" ], where "0" is none-entity. Translate to English and give the corresponding named entities. LLM response: The given Vietnamese text translates to: "[it is just like my mom and dad who were once obese and got this disease or that disease, people close to me, I know, have been]". The corresponding named entities, based on the positions provided, are: - "B-GENDER", "B-GENDER" → **"mom" and "dad"** (Gender entities) - "B-DISEASESYMTOM", "I-DISEASESYMTOM" → **"obese"** (Disease/Symptom) - "B-DISEASESYMTOM", "I-DISEASESYMTOM" → **"disease"** (Disease/Symptom) So, the translation with the named entities marked would be: "[it is just like **mom** and **dad** who were once **obese** and got this **disease** or that **disease**, people close to me, I know, have been]".
Appendix D Details about Experimental Setups
D.1 Evaluation Toolkit
We employed seqeval [^2] https://github.com/chakki-works/seqeval framework commonly used as a default evaluation framework by HuggingFace. However, this framework only works for NER on reference text. Therefore, we also employed the F1 score calculation by Shon et al. (2023) by using the SLUE toolkit [^3] https://github.com/asappresearch/slue-toolkit. This F1 score evaluates an unordered list of NE phrase and tag pairs predicted for each sentence. Our proposed modification of SLUE toolkit was also used and presented below.
D.2 Modified Evaluation of SLUE toolkit
@@
We introduce a "dummy" token strategy to replace the actual NEs. This approach upholds the focus on the classification of entities rather than the extraction of verbatim phrases, which is suitable for cases where ASR errors might skew the recognition of entities in spoken transcripts.
@@
We compute two types of overall metrics: micro and macro averages. The micro average metrics aggregate the contributions of all classes to compute the average metric, while the macro average computes per-entity type metrics and averages them, without considering the frequency of each entity type. The micro average is therefore influenced by the class distribution and will be dominated by the performance on more frequent entity types. In contrast, macro averages treat all entity types equally, providing a measure of the system’s performance across different types of NEs, regardless of their frequency in the dataset.
Appendix E NER Results by Entity Types
Tables 5-24 show the results of NER on reference text by entity types using various pre-trained language models. Tables 25-44 show the results of NER on ASR output by entity types using various pre-trained language models and ASR models.
Figure 1 shows the scatter plot of NER results on reference text by entity types using various pre-trained language models, created by Tables 5-24. Figure 2 shows the scatter plot of NER results on ASR output by entity types using various pre-trained language models, created by Tables 25-44.